OpenStudy (anonymous):

if x sin (surd) x=limit from 0 to x^2 f(t)dt, where f is a continous function, find f(4). can anybody help me with this?? thank you in advance ^^

OpenStudy (experimentx):

use latex http://www.codecogs.com/latex/eqneditor.php

OpenStudy (experimentx):

build your equation in above site, post that equation between \[ and ]\ ]\ <-- is in opposite way

OpenStudy (anonymous):

i've been doing it manually... dooh!

OpenStudy (anonymous):

owhh,,thank you for the site,,here i attach the equation build,i hope correct^^

OpenStudy (experimentx):

what is that II before x in sin??

OpenStudy (anonymous):

it's a surd or radical,,^^( i hope i choose coeerect symbol,,sorry)

OpenStudy (anonymous):

no!,,sorry!!,,it's a pie

OpenStudy (experimentx):

Ah yes ... use second fundamental theorem of calculus \[ \frac{d}{dx} (x \sin \pi x) = (f(x) - f(0)) \frac {dx^2}{dx}\]

OpenStudy (experimentx):

check on it ... i might be wrong ... haven't reviewed it for quite a while

OpenStudy (anonymous):

owhh,,thank you so much,,i'll try do do it from here,,if anything i'll ask you again..thank you so much for your effort^_^

OpenStudy (experimentx):

Oh ... sorry i was wrong!! \[\frac{d}{dx} (x \sin \pi x) = f(x) \frac {dx^2}{dx} \]

OpenStudy (anonymous):

owh,,experimentX,,here i got the solution,but do not really understand it~ sorry,the image not so clear,i snap with my phone,,

OpenStudy (experimentx):

sorry ... my last one was also wrong!! the right one should be \[ \frac{d}{dx} (x \sin \pi x) = f(x^2) \frac {dx^2}{dx} \]

OpenStudy (experimentx):

can't see your picture

OpenStudy (experimentx):

http://www.wolframalpha.com/input/?i=%28differentiate+x*sin%28pi*x%29%29+where+x%3D2 answer should be \( \frac{2pi}4 = \frac{pi}2 \)

OpenStudy (anonymous):

sorry for the inconvenience,, owhh,,really?,,thank you so much^^

OpenStudy (experimentx):

yw